Let e = (1, 0) and f = (1, 1) be two points in 

R2 . Find ix - yl, 12x - y I , where x = e + f, 

y = 2e + 2f


1
Expert's answer
2021-04-29T14:38:51-0400

"x = e + f = (1, 0)+(1, 1)= (2, 1) \\
y = 2e + 2f = 2x = (4, 2) \\
|x-y| = |x-2x| = |x| = \sqrt{2^2+1^2} =
\sqrt{5} \\
|2x-y| = |2x-2x| = |0| = 0"